# seq-JEPA: Autoregressive Predictive Learning of Invariant-Equivariant World Models

### v1

@@ 5.6 Role of action conditioning
Random noise 58.16 $\pm$ 3.2 0.20 $\pm$ 0.07
Ablated predictor act cond. 87.17 $\pm$ 0.3 0.37 $\pm$ 0.06
Ablated transformer act cond. 86.33 $\pm$ 0.1 0.53 $\pm$ 0.05
Act. embedding dim
Non-learnable act emb 85.68 $\pm$ 0.2 0.34 $\pm$ 0.04
$adim=16$ 86.29$\pm$ 0.4 0.70 $\pm$ 0.01
$adim=64$ 87.11$\pm$ 0.2 0.70 $\pm$ 0.02
$adim=128$ (default) 87.41 $\pm$ 0.5 0.71 $\pm$ 0.02
$adim=256$ 87.26$\pm$ 0.6 0.72 $\pm$ 0.00
@@
Figure 8: Effect of training and inference sequence length on seq-JEPA’s performance; left:: Equivariant performance ($R^{2}$) for different training sequence lengths (inference sequence length fixed); middle: Classification performance on top of aggregate representations versus training sequence length (inference sequence length fixed); right: Classification performance on top of aggregate representations for different inference sequence lengths (training sequence length fixed)
5.7 Role of training and inference sequence lengths
@@
In order to study the effect of training and inference sequence lengths on our framework’s performance, we trained seq-JEPA with different training sequence lengths and measured both equivariance of encoder representations and linear probe accuracy using different inference sequence lengths to form the aggregate representation (Figure 8). A general positive trend can be observed in equivariance performance for different transformations when increasing the training sequence length (Figure 8, left panel), which can be the impact of seeing more action-conditioned representations during training. For linear classification (invariance-related performance; Figure 8, middle panel), the models trained on 3DIEBench and via predictive learning across saccades benefit from a longer training sequence length. However, in the augmentation setting with CIFAR100, a small drop in accuracy is observed for higher training sequence lengths. The observed differences across evaluation tasks likely stem from specific task requirements. We hypothesize that leveraging and aggregating a sequence of action-observation pairs, i.e. seq-JEPA’s architectural inductive bias, is most effective in settings where the downstream task benefits from sequential observation. In the case of object rotations in 3DIEBench, seeing an object from multiple angles is indeed beneficial in recognizing the object’s category, which explains the improved classification accuracy with increased sequence length. Similarly, in the case of predictive learning across saccades, each eye movement and its subsequent glance provides additional information that can be leveraged for learning a richer aggregate representation. Such an advantage may not be present when using artificially-augmented images which can also explain the classification performance gap between seq-JEPA and the best-performing baselines in this setting (Table 2). Finally, during inference (Figure 8, right panel), all transformation settings benefit from more views and a richer context to improve the quality of aggregate representation.
@@ 6 Limitations and Future Perspectives
In this paper, we proposed seq-JEPA, a world modeling paradigm for learning architecturally segregated transformation-invariant and -equivariant representations. Although we have shown the feasibility of our world model for multiple transformation settings in image modality, we have not yet explored other modalities such as video, text, or multi-modal data. Given the architecture of seq-JEPA, it is possible to utilize the sequence aggregator to fuse representations from multiple modalities. Another promising direction is to build efficient world models for embodied agents with egocentric vision via scaling our approach to larger image and video datasets and take advantage of a sequence of foveated observations based on salience in a setup similar to predictive learning across saccades. Our transfer learning results on ImageNet-1k in Section B.6 hints at the possibility of building such a world model. Finally, similar to other equivariant SSL techniques, seq-JEPA requires knowledge of transformation groups. Developing a group-agnostic equivariant method is a challenge that remains largely unexplored in SSL.
Acknowledgements
@@ A.1 Data preparation
3DIEBench. The original $256\times 256$ images are resized to a $128\times 128$ resolution for all experiments. Normalization is done using the means and standard deviations provided by Garrido et al. (2023), i.e., $\mu=[0.5016,0.5037,0.5060]$ and $\sigma=[0.1030,0.0999,0.0969]$ for the three RGB channels, respectively.
@@
STL10. In order to extract the saliencies, we resize images to $512\times 512$, feed them to the pre-trained DeepGaze IIE (Linardos et al., 2021), resize the output saliencies back to $96\times 96$, and store them alongside original images. We use normalization parameters $\mu=[0.4467,0.4398,0.4066]$, $\sigma=[0.2241,0.2215,0.2239]$ for the three RGB channels, respectively. After sampling fixations from saliencies, the patches are extracted from the image to simulate foveation are $32\times 32$ (compared to the full image size of $96\times 96$).
Transformation parameters. For 3DIEBench, we use the rotation and color parameters provided with images for action conditioning as done in (Garrido et al., 2023). For the augmentation setting, we use the parameters corresponding to each of the three augmentations and form the action as the relative augmentation vector between two images. For crop, we use four variables, i.e., vertical and horizontal coordinate, and height and width. For color jitter, we use four variables: brightness, contrast, saturation, and hue. For blur, we use one variable: the standard deviation of the blurring kernel. In predictive learning across saccades, the action is a 2-d vector, i.e., the normalized relative $(x,y)$ coordinate between two patches.
@@ A.2 Training and evaluation details
For linear probing, we follow a common SSL protocol and train a linear classifier on top of frozen representations with a batch size of 256 for 300 epochs using the Adam optimizer with default hyperparameters. For action prediction, we follow a similar protocol as SIE (Garrido et al., 2023). Specifically, for rotation, color jitter, and crop, we train an MLP regressor with a hidden dimension of 1024 and ReLU activation for 300 epochs. For color (in 3DIEBench), blur (in the augmentation setting), and position (in predictive learning across saccades), we use a linear regressor and train it for 50 epochs. For path integration experiments, the same regressor architectures as the equivariance evaluation heads are used, i.e., the MLP for angular path integration and the linear regressor for saccade path integration. All regression heads are trained using the Adam optimizer with default hyperparameters.
@@ A.3 Architectural details
We use a a projection head with 2048-2048-2048 intermediate dimensions. The predictor has a hidden dimension of 512-d with ReLU activation. We use the same EMA setup outlined in the original paper (Grill et al., 2020), i.e., the EMA parameter $\tau$ starts from $\tau_{\text{base}}=0.996$ and is increased following a cosine schedule.
@@ B.1 Evaluation results on 3DIEBench for models conditioned on rotation and color
We further provide evaluation results for seq-JEPA and equivariant baselines on 3DIEBench conditioned on both rotation and color transformations in Table 5. The results show that all methods suffer a drop in classification performance and become highly sensitive to color. Similar performance degradations have been previously observed with 3DIEBench (Garrido et al., 2023; Gupta et al., 2024), yet without a convincing explanation. One possible explanation in line with Principle II in Wang et al. (2024) is that in 3DIEBench, floor and light hue are fairly decoupled from class information (low class relevance). Therefore, forcing the encoder to encode color information would cause class information to be lost, resulting in degradation of classification accuracy.
Table 5: Quantitative evaluation of learned representations on 3DIEBench for rotation and color prediction (equivariance) and linear probe classification (invariance). All models are conditioned on both rotation and color.
@@ B.2 Measuring invariance of aggregate representations
In order to have a sense of invariance of the aggregate representation, we tried to predict the individual rotation parameters of each view for a seq-JEPA model with $M_{val}=3$ trained on 3DIEBench and conditioned on rotation. To avoid leaking the rotation parameters to $z_{AGG}$ during prediction, we zero out the action embeddings. The mean $R^{2}$ score obtained by predicting individual rotations of the three views using $z_{AGG}$ is $0.141\pm 0.002$; which indicates that the aggregate representation has not encoded rotation in any of the views.
@@ B.3 Complete evaluation results for linear probing on top of aggregate representations
In this section, we provide our complete evaluation results for linear probing on top of seq-JEPA’s aggregate representations for three transformation sets with different training and inference sequence lengths. Figure 9 shows the top-1 accuracy on 3DIEBench models conditioned on rotation. Figure 10 shows top-1 accuracy on STL-10 for models trained via predictive learning across saccades. Figures 11 and 12 show top-1 accuracy on CIFAR100 and Tiny ImageNet, respectively, with different types of action conditioning (crop, color jitter, blur, or all three). We can see the general trends of Figure 8 discussed in Section 5.7 also in these heatmaps.
@@ B.4 Comparison of evaluation results on encoder representations and aggregate representations
For completeness, we provide linear probe classification on encoder representations for different transformation settings in Table 6 and compare them with accuracy on aggregate representations for different inference evaluation lengths. The aggregate representation generally achieves a higher classification performance thanks to the architectural inductive bias in seq-JEPA.
@@
STL-10 None 4 61.38 62.21 69.06 71.12
STL-10 Position 4 81.20 71.45 81.53 83.67
STL-10 Position (no saliency) 4 79.29 63.14 76.93 80.09
STL-10 Position (no IoR) 4 72.49 68.95 76.84 78.21
B.5 Visualization of UMAP projections for models without action conditioning
@@
In Figure 13 we provide the UMAP projections for seq-JEPA representations trained on 3DIEBench but without action conditioning. Figure 14 shows UMAP projections for seq-JEPA representations trained on STL-10 via predictive learning across saccades without action conditioning. In these visualizations, the color gradients that implied high equivariance in UMAPs of action-conditioned models are either absent or less pronounced.
@@
Figure 13: Visualization of 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations. The model is trained on 3DIEBench without any action conditioning with $M_{tr}=3$ and $M_{val}=5$. Encoder representations for each view observation, color-coded by class category (left) and rotation angle (middle). The aggregate token representation for $M_{val}=5$ observations during inference, color-coded by class category (right).
@@
Figure 14: Visualization of 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations. The model is trained on STL-10 via predictive learning across saccades with no action conditining and with $M_{tr,val}=4$. Encoder representations for each view observation, color-coded by class category (left) and $X$ and $Y$ pixel coordinates (middle panels). The aggregate token representation for $M_{val}=4$ observation patches during inference, color-coded by class category (right).
B.6 Transfer learning results on ImageNet-1k
In this section, we consider transfer learning on ImageNet-1k using our frozen model trained on STL-10 via predictive learning across saccades. We follow the same evaluation protocol as the in-distribution setting and train a linear probe on ImageNet-1k training set and evaluate it on the validation set using the aggregate representation for different inference sequence lengths. We use DeepGaze-IIE to extract saliency maps of ImageNet-1k, resize them to $224\times 224$, and use them to extract the foveated patches. Figure 15 shows the top-1 accuracy on ImageNet-1k for two different patch sizes (32 and 84). These results hint at the possibility of scaling seq-JEPA to larger data regimes by striking a trade-off between the patch size and sequence length. For example, using full-resolution $224\times 224$ images, a forward pass through a ResNet-18 encoder takes approximately $1.83\times 10^{9}$ flops, while a forward pass of seq-JEPA with the default transformer encoder, $M_{val}=4$ and $84\times 84$ patches takes $1.25\times 10^{9}$ flops. Using the CIFAR variant of ResNet-18 with $32\times 32$ patches and $M_{val}=8$ requires $2.35\times 10^{9}$ flops.

### v2

@@ 4.5 Ablations: Action Conditioning Is Key
No predictor cond. 87.17 $\pm$ 0.3 0.37 $\pm$ 0.06
No transformer cond. 86.33 $\pm$ 0.1 0.53 $\pm$ 0.05
Act. embedding dim.
Non-learnable 85.68 $\pm$ 0.2 0.34 $\pm$ 0.04
$a_{\text{dim}}=16$ 86.29 $\pm$ 0.4 0.70 $\pm$ 0.01
$a_{\text{dim}}=64$ 87.11 $\pm$ 0.2 0.70 $\pm$ 0.02
$a_{\text{dim}}=128$ (default) 87.41 $\pm$ 0.5 0.71 $\pm$ 0.02
$a_{\text{dim}}=256$ 87.26 $\pm$ 0.6 0.72 $\pm$ 0.00
@@
Figure 7: Effect of training and inference sequence length on seq-JEPA’s performance; left:: Equivariant performance ($R^{2}$) versus training sequence length; middle: Classification accuracy versus training sequence length; right: Classification accuracy versus inference sequence length.
4.6 Scaling Properties: Role of Training and Inference Sequence Lengths
@@
We study the effect of both training and inference sequence lengths on performance across tasks, i.e., scalability of seq-JEPA in terms of context length (Figure 7). We draw inspiration from recent findings in large-scale foundation models, where increased training and inference context—whether in text (Brown et al., 2020; Touvron et al., 2023), vision (Zellers et al., 2021; Chen et al., 2021), or video (Bain et al., 2021; Arnab et al., 2021)—consistently leads to stronger representations. We observe that equivariance generally improves with longer training sequences (left panel), likely due to increased exposure to action-conditioned transitions. Classification performance on 3DIEBench and STL-10 (middle panel) similarly benefits. In contrast, on CIFAR100 and Tiny ImageNet with synthetic augmentations, longer training sequences slightly decrease classification performance. We hypothesize that leveraging and aggregating a sequence of action-observation pairs, i.e. seq-JEPA’s architectural inductive bias, is most effective in settings where the downstream task benefits from sequential observations. In the case of object rotations in 3DIEBench, seeing an object from multiple angles is indeed beneficial in recognizing the object’s category, which explains the improved classification accuracy with increased training sequence length. Similarly, in the case of predictive learning across saccades, each eye movement and its subsequent glance provides additional information that can be leveraged for learning a richer aggregate representation.
At inference time, all datasets benefit from longer context lengths ($M_{val}$), confirming that richer aggregate representations yield stronger performance (right plot). This scalability via sequence lengths opens avenues for efficient representation learning with small foveated patches in lieu of full-frame inputs, mirroring how foundation models scale with input tokens at test time. Together with our transfer learning results on ImageNet-1k (Appendix B.7), these findings suggest that seq-JEPA’s architectural inductive bias enables graceful scaling via longer sequence lengths.
5 Related Work
Non-Generative World Models and Joint-Embedding Predictive Architectures. Non-generative world models predict the effect of transformations or actions directly in latent space, avoiding reconstruction in pixel space. This includes contrastive SSL methods that model transformed views from context representations (van den Oord et al., 2019; Gupta et al., 2024), as well as approaches in model-based reinforcement learning (RL) to improve sample efficiency, generate intrinsic rewards, or capture environment transitions (Schwarzer et al., 2021; Khetarpal et al., 2025; Ni et al., 2024; Tang et al., 2023; Guo et al., 2022). Joint-embedding predictive architectures (LeCun, 2022) form a structured subclass of non-generative world models. Rather than aligning or contrasting representations directly, they introduce an asymmetric predictor conditioned on transformation parameters to infer the outcome of an action applied to a latent view. Examples include I-JEPA(Assran et al., 2023), which predicts masked regions from positional cues, and IWM (Garrido et al., 2024), which conditions on augmentation parameters. JEPAs have been recently extended to physical reasoning (Garrido et al., 2025) and offline planning (Sobal et al., 2025), illustrating the framework’s versatility in structured predictive representation learning.
Equivariant SSL. Equivariant SSL methods aim to retain transformation-specific information in the latent space, typically by augmenting invariant objectives with an additional equivariance term. Some approaches directly predict transformation parameters (Lee et al., 2021; Scherr et al., 2022; Gidaris et al., 2018; Gupta et al., 2024; Dangovski et al., 2022). Methods such as EquiMod (Devillers and Lefort, 2022) and SIE (Garrido et al., 2023) predict the effect of a transformation in latent space via a predictor in addition to their invariant objective. SEN (Park et al., 2022) similarly predicts transformed representations but omits the invariance term. Xiao et al. (2021) use contrastive learning with separate projection heads for each augmentation, treating same-augmentation pairs as negatives. ContextSSL (Gupta et al., 2024) conditions representations on both current actions and recent context using a memory module and employs a dual predictor for transformation prediction to avoid collapsing to invariance. Other approaches (Shakerinava et al., 2022; Gupta et al., 2023; Yerxa et al., 2024) do not require explicit transformation parameters but instead enforce equivariance by applying the same transformation to multiple view pairs and minimizing a distance-based loss. Finally, action-conditioned JEPAs like IWM incorporate augmentation parameters and mask positions by conditioning the predictor to induce equivariance without additional objectives, though they still face the invariance-equivariance tradeoff depending on predictor capacity (Garrido et al., 2024).
Positioning of Our Work. seq-JEPA belongs to the family of joint-embedding predictive architectures and is a non-generative world model. Unlike most equivariant SSL methods, seq-JEPA does not rely on an equivariance loss or transformation prediction objective, nor does it require view pairs with matched transformations. Instead, it introduces architectural inductive biases to JEPAs (Assran et al., 2023; Garrido et al., 2024) to learn separate invariant and equivariant representations, which emerge naturally from its action-conditioned sequential predictive learning. Furthermore, in contrast to ContextSSL, which extends the two-view contrastive setting of contrastive predictive coding van den Oord et al. (2019) using a transformer decoder projector conditioned on previous views via key-value caching, seq-JEPA operates on sequences of action-observation pairs in an online end-to-end manner by incorporating a sequence model (e.g. a transformer encoder) as a learned working memory during both training and inference. Moreover, while ContextSSL aims to adapt equivariance to recent transformations by dynamically modifying the training distribution, seq-JEPA is designed to explicitly learn both invariant and equivariant representations with respect to a specified set of transformations, and is also well-suited for downstream tasks that require multi-step observation aggregation (e.g., predictive learning across saccades and path integration).
@@ 6 Limitations and Future Perspectives
We have validated viability of seq-JEPA across a range of transformations in the image domain. Here, we discuss limitations and possible future directions. First, the transformer-based aggregator in seq-JEPA could support multi-modal fusion across language, audio, or proprioceptive inputs—enabling multi-modal world modeling and generalization. Second, while our results show clear benefits from longer sequences, we have experimented with relatively short training and inference sequence lengths as the ResNet backbone is also trained end-to-end. The sequence scaling trends observed in the paper suggest that seq-JEPA can benefit from longer context windows over representations of pre-trained backbones as in (Pang et al., 2023; Lin et al., 2024). Third, our method assumes access to a known transformation group (e.g., $SO(3)$ for 3D rotations). Designing group-agnostic or learned transformation models (Finzi et al., 2021) without access to transformation parameters or pairs of same transformation is an open challenge in equivariant SSL that future work may tackle. Finally, our preliminary ImageNet-1k transfer results (Appendix B.7) point to potential for broader generalization. Scaling seq-JEPA to larger foveated image settings or video and multi-modal datasets such as Song et al. (2023) could support the development of lightweight, saliency-driven agents capable of learning efficiently from partial observations in embodied settings with a limited field of vision.
Acknowledgments and Disclosure of Funding
@@ A.1 Data preparation
3DIEBench. The original $256\times 256$ images are resized to a $128\times 128$ resolution for all experiments. Normalization is done using the means and standard deviations in Garrido et al. [2023], i.e., $\mu=[0.5016,0.5037,0.5060]$ and $\sigma=[0.1030,0.0999,0.0969]$ for the three RGB channels, respectively.
@@
STL10. In order to extract the saliencies, we resize images to $512\times 512$, feed them to the pre-trained DeepGaze IIE [Linardos et al., 2021], resize the output saliencies back to $96\times 96$, and store them alongside original images. We use normalization parameters $\mu[0.4467,0.4398,0.4066]$, $\sigma=[0.2241,0.2215,0.2239]$ for the three RGB channels, respectively. After sampling fixations from saliencies, the patches that are extracted from the image to simulate foveation are $32\times 32$ (compared to the full image size of $96\times 96$). For IoR, we zero-out a circular area with radius of $16$ around each previous fixation.
Transformation parameters. For 3DIEBench, we use the rotation and color parameters provided with images for action conditioning as done in Garrido et al. [2023]. For the augmentation setting, we use the parameters corresponding to each of the three augmentations and form the action as the relative augmentation vector between two images. For crop, we use four variables, i.e., vertical and horizontal coordinate, and height and width. For color jitter, we use four variables: brightness, contrast, saturation, and hue. For blur, we use one variable: the standard deviation of the blurring kernel. In predictive learning across saccades, the action is a 2-d vector, i.e., the normalized relative $(x,y)$ coordinate between two patches.
@@ A.2 Training and evaluation details
Additional Training Details.
We used the PyTorch framework for training all models. For experiments that use CIFAR100 and low-resolution STL-10 patches in predictive learning across saccades, we use the CIFAR variant of ResNet-18. For models trained with AdamW, we used with default $\beta_{1}$ and $\beta_{2}$, a weight decay of $0.001$, and a learning rate of $4\times 10^{-4}$ with a linear warmup for $20$ epochs starting from $10^{-5}$ followed by a cosine decay back to $10^{-5}$. For models trained with Adam, we use the Adam optimizer with a learning rate of $10^{-3}$, default $\beta_{1}$ and $\beta_{2}$, and no weight decay.
@@
For linear probing, we follow a common SSL protocol and train a linear classifier on top of frozen representations with a batch size of 256 for 300 epochs using the Adam optimizer with default hyperparameters. For action prediction, we follow a similar protocol as SIE [Garrido et al., 2023]. Specifically, for rotation, color jitter, and crop, we train an MLP regressor with a hidden dimension of 1024 and ReLU activation for 300 epochs. For color (in 3DIEBench), blur (in the augmentation setting), and position (in predictive learning across saccades), we use a linear regressor and train it for 50 epochs. For path integration experiments, the same regressor architectures as the equivariance evaluation heads are used, i.e., the MLP for angular path integration and the linear regressor for saccade path integration. All regression heads are trained using the Adam optimizer with default hyperparameters.
@@ A.3 Architectural details
We use a a projection head with 2048-2048-2048 intermediate dimensions. The predictor has a hidden dimension of 512-d with ReLU activation. We use the same EMA setup outlined in the original paper [Guo et al., 2020], i.e., the EMA parameter $\tau$ starts from $\tau_{\text{base}}=0.996$ and is increased following a cosine schedule.
@@ B.1 Evaluation results on 3DIEBench for models conditioned on rotation and color
Table 5 reports performance of seq-JEPA and several equivariant baselines when conditioned on both rotation and color in 3DIEBench. All methods suffer a drop in classification performance and become highly sensitive to color. Similar performance degradations have been previously observed with 3DIEBench [Garrido et al., 2023, Gupta et al., 2024], though without a clear explanation. One plausible explanation aligned with Principle II in Wang et al. [2024], is that color in 3DIEBench, i.e., floor and light hue, are weakly correlated with class labels in 3DIEBench (low class relevance). Therefore, forcing the encoder to encode color information would cause class information to be lost, resulting in degradation of classification accuracy.
Table 5: Evaluation on 3DIEBench for rotation and color prediction (equivariance) and linear probe classification (invariance). All models are conditioned on both rotation and color.
@@ B.2 Measuring invariance of aggregate representations
In order to have a more direct measure of invariance of the aggregate representation, we trained a model on 3DIEBench with $M_{val}=3$ and rotation conditioning, then attempted to predict the individual rotation parameters of each input view from $z_{\text{AGG}}$. To prevent leakage of rotation information during prediction, all action embeddings were zeroed out. The resulting mean $R^{2}$ score obtained by predicting individual rotations of the three views using $z_{AGG}$ is $0.141\pm 0.002$; indicating that the aggregate representation contains little to no rotation-specific information.
@@ B.3 Complete evaluation results for linear probing on top of aggregate representations
We provide our complete evaluation results for linear probing on top of seq-JEPA’s aggregate representations for our three transformation settings with different training and inference sequence lengths. Figure 8 shows the top-1 accuracy on 3DIEBench models conditioned on rotation. Figure 9 shows top-1 accuracy on STL-10 for models trained via predictive learning across saccades. Figures 10 and 11 show top-1 accuracy on CIFAR100 and Tiny ImageNet, respectively, with different types of action conditioning (crop, color jitter, blur, or all three). These heatmaps reflect the same trends observed in Figure 7 and discussed in Section 4.6, illustrating the consistent effect of sequence length on representation quality.
@@ B.4 Comparison of evaluation results on encoder representations and aggregate representations
For completeness, we provide linear probe classification on encoder representations for different transformation settings in Table 6 and compare them with accuracy on aggregate representations for different inference evaluation lengths. The aggregate representation generally achieves a much higher classification performance thanks to the architectural inductive bias in seq-JEPA.
@@
STL-10 None 4 61.38 62.21 69.06 70.45
STL-10 Position 4 81.20 71.45 81.53 83.44
STL-10 Position (no saliency) 4 79.29 63.14 76.93 79.85
STL-10 Position (no IoR) 4 72.49 68.95 76.84 77.97
B.5 Additional UMAP Visualizations
@@
In Figure 12, we visualize the UMAP projections of seq-JEPA representations trained on 3DIEBench without action conditioning. Similarly, Figure 14 shows projections for models trained on STL-10 via predictive learning across saccades, also without action conditioning. Compared to the action-conditioned counterparts (Figures 5 and 13), these projections exhibit weaker or no smooth color gradients in their corresponding transformation-colored UMAP—indicating reduced equivariance to transformation parameters.
@@
Figure 12: 2-D UMAP projections of seq-JEPA representations on 3DIEBench without action conditioning ($M_{tr}=3$, $M_{val}=5$). Encoder outputs are color-coded by class (left) and rotation angle (middle); aggregate token representations are color-coded by class (right).
@@
Figure 13: 2-D UMAP projections of seq-JEPA representations on STL-10 with action conditioning ($M_{tr}=M_{val}=4$). Encoder outputs are color-coded by class (top-left) and by $X$/$Y$ fixation coordinates (bottom); the aggregate token is color-coded by class (top-right).
Figure 14: 2-D UMAP projections of seq-JEPA representations on STL-10 without action conditioning ($M_{tr}=M_{val}=4$). Encoder outputs are color-coded by class (left) and fixation coordinates (middle); the aggregate token is color-coded by class (right).
B.6 Details of Path Integration Experiments.
While an agent executes a sequence of actions in an environment, transitioning from an initial state to a final state, it should be capable of tracking its position by integrating its own actions. This is also a crucial cognitive ability that enables animals to estimate their current state in their habitat [McNaughton et al., 2006]. Here, we evaluate whether seq-JEPA is capable of path integration. Given the sequence of observations $\{x_{i}\}_{i=1}^{M+1}$ generated from transformations $\{t_{i}\}_{i=1}^{M+1}$ and the corresponding relative action embeddings $\{a_{i}\}_{i=1}^{M}$, we define the task of path integration over the sequence of actions as predicting the relative action that would directly transform $x_{1}$ to $x_{M+1}$ given $z_{AGG}$ and $a_{M}$. In other words, given the aggregate representation of a sequence of action-observation pairs and the next action, we would like to predict the overall position change from the starting point ($x_{1}$) to the end point ($x_{M+1}$). We consider path integration for rotation angles in 3DIEBench and across eye movements with STL-10. For rotations, the task is integrating a series of object rotations from the first view to the last, i.e. angular path integration. For eye movements, the task is integrating the eye movements from the first saccade to the last, i.e. visual path integration. To measure path integration performance for inference sequence length $M$, we train a regression head on top of the concatenation of $z_{AGG}$ and $a_{M}$ to predict the transformation from $x_{1}$ to $x_{M+1}$. Figure 6 shows that seq-JEPA performs well in both angular and visual path integration. The red curve corresponds to the performance of the original seq-JEPA. The blue curve corresponds to experiments in which the action embeddings are ablated (zeroed-out during inference for all views). The green curve corresponds to experiments in which the encoder (visual) representations are ablated during inference. As expected, path integration becomes more difficult as the number of observations increases (red curves). Ablating action conditioning (blue curves) results in failure of path integration. On the other hand, ablating the visual representations (green curves) results only in a small performance drop compared to the original model, indicating that action conditioning is the key factor that enables path integration.
B.7 Transfer learning results on ImageNet-1k
To evaluate generalization of our model beyond STL-10, we assess transfer performance of the model trained on STL-10 patches via predictive learning across saccades on ImageNet-1k. We follow the same linear probing protocol as in-distribution evaluations: we freeze the ResNet and transformer encoder and train a linear classifier on aggregate representations generated from foveated patches.
We extract saliency maps for ImageNet-1k images using DeepGaze IIE, resize them to $224\times 224$, and sample sequences of patches sized $32\times 32$ or $84\times 84$ for this out-of-distribution (OOD) evaluation setting. Figure 15 shows top-1 linear probe accuracy on ImageNet-1k validation set across varying inference sequence lengths ($M_{val}$). For both patch sizes, performance improves with longer inference sequences, validating seq-JEPA’s ability to benefit from extended context even in this difficult OOD ImageNet-1k setting. These results echo findings in the main experiments for different training and inference sequence lengths and confirming the possibility of model’s scalability in terms of data, parameter count, and compute.